# A Low-Overhead Incorporation-Extrapolation based Few-Shot CSI Feedback Framework for Massive MIMO Systems

### v1

@@ V-A2 Experimental Settings
The number of feedback bits $B$ is set to $64$, $128$, $256$, $512$, or $1024$ for comprehensive performance comparison. For a fair comparison, all methods exploit $2$-bit uniform quantization. General hyper-parameter settings for the proposed IEFSF are shown in Table II.
@@ V-B Performance Evaluation of Proposed Data Augmentation Methods for Few-Shot CSI Feedback
The performance of the proposed knowledge-driven data augmentation method and the AIGC-based data augmentation method under indoor and outdoor scenarios is shown in Fig. 6 and Fig. 6. In these figures, the horizontal axis represents the summation of training and validation samples, while the vertical axis represents the average GCS $\bar{\rho}$. For readability, these two data augmentation methods are denoted with ‘KDDA’ and ‘EGAN’ in the figures. And the feedback bits $B=256$. As depicted in Fig. 6 and Fig. 6, the proposed methods show comparable performance in augmenting few-shot CSI feedback with only $100$ $500$ collected samples for the indoor scenario and only $200$ $1,000$ collected samples for the outdoor scenario. The knowledge-driven data augmentation method is applicable for scenarios with very few collected samples (e.g., only $100$-$200$ samples) while the AIGC-based method exhibits better performance when there are several hundred samples. Additionally, from Fig. 6 and Fig. 6, by combining the two data augmentation methods (denoted by ‘KDDA+EGAN’), the proposed framework can achieve a high CSI feedback accuracy with only several hundred collected samples.
To further validate the effectiveness of the proposed AIGC-based data augmentation method, we compare it with the existing state-of-the-art (SOTA) AIGC-based data augmentation method for CSI feedback, i.e., the ChannelGAN-based method[30]. For fare comparison, $20,000$ samples and $40,000$ samples are generated with the ChannelGAN for the indoor scenario and outdoor scenario, respectively. From Fig. 6, Fig. 6, and Table III, the EGAN-based method outperforms the ChannelGAN-based method given various numbers of training and validation samples under the scenarios, and the EGAN is much more efficient than the ChannelGAN. This is mainly due to the following three reasons: 1) In comparison to the ChannelGAN which learns to model the channel, the proposed EGAN directly learns to generate eigenvector-based CSI matrices, making it much more computationally efficient; 2) Thanks to the proposed two-scale generator and discriminator architecture, the EGAN learns both the distribution of full-dimensional CSI matrices and the distribution of low-dimensional CSI matrices, leading to more accurate distribution learning from limited collected samples; 3) With the assistance of the knowledge-driven data augmentation method, the EGAN can be trained with a high generalization capability.
Figure 5: Average GCS of the proposed data augmentation methods and the ChannelGAN-based method versus the numbers of samples for training and validation under the indoor scenario.
Figure 5: Average GCS of the proposed data augmentation methods and the ChannelGAN-based method versus the numbers of samples for training and validation under the indoor scenario.
Figure 6: Average GCS of the proposed data augmentation methods and the ChannelGAN-based method versus the numbers of samples for training and validation under the outdoor scenario.
@@
TABLE III: Computational complexity comparison of the proposed EGAN and the ChannelGAN.
@@
V-C Performance Comparison Among Proposed IEFSF and Baselines
The proposed IEFSF is compared with the following baselines: 1) CsiNet+[9], 2) bi-ImCsiNet[15] 3) TransNet[11], and 4) SampleDL[13]. For fair comparison, all baselines are adapted for eigenvector-based CSI feedback. In the proposed IEFSF, the ”KDDA+EGAN” data augmentation method is adopted.
Figure 7: Performance comparison of the proposed IEFSF and baselines versus the numbers of samples for training and validation under the indoor scenario.
Figure 7: Performance comparison of the proposed IEFSF and baselines versus the numbers of samples for training and validation under the indoor scenario.
Figure 8: Performance comparison of the proposed IEFSF and baselines versus the numbers of samples for training and validation under the outdoor scenario.
We first evaluate the impact of the number of training and validation samples. As depicted in Fig. 8 and Fig. 8, the IEFSF and the SampleDL show the best and the worst CSI feedback performance under both indoor and outdoor scenarios given few collected samples, respectively. Specifically, under the system configuration of $32$ antennas and $100$ MHz bandwidth with $1024$ subcarriers, the IEFSF can achieve up to $0.94$ average GCS with only $500$ collected samples for the indoor scenario and up to $0.85$ average GCS with only $1,000$ collected samples for the outdoor scenario. This demonstrates the superiority of the proposed IEFSF for few-shot CSI feedback.
@@ V-C Performance Comparison Among Proposed IEFSF and Baselines
The impact of feedback bits on the CSI feedback performance is also investigated. It can be observed from Fig. 10 and Fig. 10 that more feedback bits bring about higher feedback accuracy. However, as the number of feedback bits increases to a relatively large value (e.g., $256$), the performance improvement becomes limited, showing a marginal effect. Besides, increasing the number of feedback bits will also lead to increasing computational complexity. Therefore, a proper setting of feedback bits deserves a dedicated selection in practical applications. In addition, as shown in Fig. 10 and Fig. 10, the IEFSF respectively achieves around $0.93$ and $0.84$ average GCS under the indoor and outdoor scenarios with only $64$ feedback bits, while the baselines merely achieve at most $0.78$ and $0.55$ average GCS under both scenarios with $1024$ feedback bits. This indicates the proposed IEFSF can achieve high CSI feedback accuracy with much low feedback overhead (e.g., $16$ times lower feedback overhead than the baselines). It is worth mentioning that performance degradation of SampleDL can be observed when the number of feedback bits increases from $256$ onwards. This is because the SampleDL adopts a two-step training scheme, i.e., training the CSI feedback network first and then training the interpolating network. Given a large number of feedback bits while training the CSI feedback network with few training samples, the trained CSI feedback network is prone to overfitting to the first-stage CSI feedback. As a consequence, the interpolating network can not be trained properly from biased outputs from the trained CSI feedback network. Different from the SampleDL, the proposed IEFSF adopts joint training of CSI feedback and interpolation, leading to a stable feedback accuracy improvement with the increase of feedback bits.
@@
We further compare the computational complexity of the proposed IEFSF with baselines. The computational complexity of all schemes with a fixed setting of feedback bits $B=256$ is shown in Table IV. From Table IV, the IEFSF and the SampleDL show computational efficiency over common DL-based CSI feedback schemes, i.e., CsiNet+, bi-ImCsiNet, and TransNet. This is because the SampleDL and the proposed IEFSF greatly reduce the CSI dimension for feedback before the CSI compression, such that the computational complexity can be significantly reduced. Further comparing the proposed IEFSF with the SampleDL, the proposed IEFSF has a significantly smaller time complexity than the SampleDL (i.e., $25.1$ times smaller) at the cost of a slightly larger space complexity (i.e., $1.3$ times larger than the SampleDL).
TABLE IV: Computational complexity of the proposed IEFSF and baselines.

### v2

@@ V-A2 Experimental Settings
The number of feedback bits $B$ is set to $64$, $128$, $256$, $512$, or $1024$ for comprehensive performance comparison. For a fair comparison, all methods exploit $2$-bit uniform quantization. General hyper-parameter settings for the proposed IEFSF are shown in Table II. Specifically, we set $\lambda_{1}=\lambda_{2}=10$ following [50] for training the EGAN.
@@
$\lambda_{1}$ $10$
$\lambda_{2}$ $10$
TABLE III: Quantified frequency domain correlations among adjacent subcarrier groups with different group granularities.
Granularity $\overline{\tilde{\rho}}$ (Indoor) $\overline{\tilde{\rho}}$ (Outdoor)
1 1.0000 0.9996
2 1.0000 0.9985
4 1.0000 0.9948
8 0.9999 0.9837
16 0.9995 0.9545
32 0.9980 0.9210
64 0.9933 0.9140
128 0.9845 0.8878
256 0.9864 0.8282
@@ V-B Performance Evaluation of Proposed Data Augmentation Methods for Few-Shot CSI Feedback
The performance of the proposed knowledge-driven data augmentation method and the AIGC-based data augmentation method under indoor and outdoor scenarios is shown in Fig. 7 and Fig. 7. In these figures, the horizontal axis represents the summation of training and validation samples (i.e., the number of collected samples), while the vertical axis represents the average GCS $\bar{\rho}$. For readability, these two data augmentation methods are denoted with ‘KDDA’ and ‘EGAN’ in the figures. And the feedback bits $B=256$. As depicted in Fig. 7 and Fig. 7, the proposed methods show comparable performance in augmenting few-shot CSI feedback with only $100$ $500$ collected samples for the indoor scenario and only $200$ $1,000$ collected samples for the outdoor scenario. The knowledge-driven data augmentation method is applicable for scenarios with very few collected samples (e.g., only $100$-$200$ samples) while the AIGC-based method exhibits better performance when there are several hundred samples. Additionally, from Fig. 7 and Fig. 7, by combining the two data augmentation methods (denoted by ‘KDDA+EGAN’), the proposed framework can achieve a high CSI feedback accuracy with only several hundred collected samples.
To further validate the effectiveness of the proposed AIGC-based data augmentation method, we compare it with the existing state-of-the-art (SOTA) AIGC-based data augmentation method for CSI feedback, i.e., the ChannelGAN-based method[33]. For fare comparison, $20,000$ samples and $40,000$ samples are generated with the ChannelGAN for the indoor scenario and outdoor scenario, respectively. From Fig. 7, Fig. 7, and Table IV, the EGAN-based method outperforms the ChannelGAN-based method given various numbers of collected samples under the scenarios, and the EGAN is much more efficient than the ChannelGAN. This is mainly due to the following three reasons: 1) In comparison to the ChannelGAN which learns to model the channel, the proposed EGAN directly learns to generate eigenvector-based CSI matrices, making it much more computationally efficient; 2) Thanks to the proposed two-scale generator and discriminator architecture, the EGAN learns both the distribution of full-dimensional CSI matrices and the distribution of low-dimensional CSI matrices, leading to more accurate distribution learning from limited collected samples; 3) With the assistance of the knowledge-driven data augmentation method, the EGAN can be trained with a high generalization capability.
Figure 6: Average GCS of the proposed data augmentation methods and the ChannelGAN-based method versus the numbers of collected samples under the indoor scenario.
Figure 7: Average GCS of the proposed data augmentation methods and the ChannelGAN-based method versus the numbers of collected samples under the outdoor scenario.
@@
TABLE IV: Computational complexity comparison of the proposed EGAN and the ChannelGAN.
@@
V-C Performance Comparison Among Proposed IEFSF and Baselines Under Few-Shot Scenarios
The proposed IEFSF is compared with the following learning-based baselines: 1) CsiNet+[9], 2) bi-ImCsiNet[15] 3) TransNet[11], and 4) SampleDL[13] under few-shot scenarios. The proposed IEFSF adopts the “KDDA+EGAN” data augmentation to tackle the challenge of lacking collected samples, while the baselines are trained with originally collected samples. In addition, we further compare the proposed IEFSF with conventional non-learning-based CSI feedback methods, including the random vector quantization (RVQ) based CSI feedback method [53] and the 5G NR Type I codebook-based CSI feedback method [14]. Note that since the RVQ-based method and the Type I codebook-based method do not learn from CSI samples, their performance does not change with the number of collected samples. For fair comparisons, all baselines are adapted for eigenvector-based CSI feedback.
Figure 8: Performance comparison of the proposed IEFSF and baselines versus the numbers of collected samples under the indoor scenario.
Figure 9: Performance comparison of the proposed IEFSF and baselines versus the numbers of collected samples under the outdoor scenario.
We first evaluate the impact of the number of collected samples with a fixed number of feedback bits $B=256$. As depicted in Fig. 9 and Fig. 9, the IEFSF and the SampleDL show the best and the worst CSI feedback performance under both indoor and outdoor scenarios given few collected samples, respectively. Specifically, under the system configuration of $32$ antennas and $100$ MHz bandwidth with $1024$ subcarriers, the IEFSF can achieve up to $0.94$ average GCS with only $500$ collected samples for the indoor scenario and up to $0.85$ average GCS with only $1,000$ collected samples for the outdoor scenario. This demonstrates the superiority of the proposed IEFSF for few-shot CSI feedback. Moreover, it can be seen from Fig. 9 that the proposed IEFSF outperforms the RVQ-based CSI feedback method with only $200$ collected samples and outperforms the Type I codebook-based CSI feedback method with only $300$ collected samples under the indoor scenario. From Fig. 9, the proposed IEFSF outperforms the RVQ-based CSI feedback method with only $200$ collected samples and outperforms the Type I codebook-based CSI feedback method with only $600$ collected samples under the outdoor scenario. These results indicate that only a few collected samples are required by the proposed IEFSF to surpass non-learning baselines and further demonstrate the practicality of our proposed IEFSF for adapting to new scenarios.
@@ V-C Performance Comparison Among Proposed IEFSF and Baselines Under Few-Shot Scenarios
The impact of feedback bits on the CSI feedback performance is also investigated. The number of collected samples is fixed at $500$ and $1000$ for the indoor scenario and the outdoor scenario, respectively in these experiments. It can be observed from Fig. 11 and Fig. 11 that more feedback bits bring about higher feedback accuracy. However, as the number of feedback bits increases to a relatively large value (e.g., $256$), the performance improvement becomes limited, showing a marginal effect. Besides, increasing the number of feedback bits will also lead to increasing computational complexity. Therefore, a proper setting of feedback bits deserves a dedicated selection in practical applications. In addition, as shown in Fig. 11 and Fig. 11, the IEFSF respectively achieves around $0.88$ and $0.77$ average GCS under the indoor and outdoor scenarios with only $16$ feedback bits. As a comparison, the learning-based baselines merely achieve at most $0.78$ and $0.55$ average GCS and the conventional non-learning-based baselines merely achieve at most $0.55$ and $0.54$ average GCS under both scenarios with $1024$ feedback bits. This indicates the proposed IEFSF can achieve high CSI feedback accuracy with much lower feedback overhead (i.e., up to $64$ times lower feedback overhead than the baselines). It is worth mentioning that performance degradation of SampleDL can be observed when the number of feedback bits increases from $256$ onwards. This is because the SampleDL adopts a two-step training scheme, i.e., training the CSI feedback network first and then training the interpolating network. Given a large number of feedback bits while training the CSI feedback network with few training samples, the trained CSI feedback network is prone to overfitting to the first-stage CSI feedback. As a consequence, the interpolating network can not be trained properly from biased outputs from the trained CSI feedback network. Different from the SampleDL, the proposed IEFSF adopts joint training of CSI feedback and interpolation, leading to a stable feedback accuracy improvement with the increase of feedback bits.
We further investigate the performance of the proposed IEFSF and baselines versus various SNRs. The number of feedback bits is fixed to $B=256$ and the number of collected samples is fixed at $500$ and $1000$ for the indoor scenario and the outdoor scenario, respectively. As shown in Fig. 13 and Fig. 13, the proposed IEFSF outperforms the baselines either with a low SNR or with a high SNR under both scenarios and higher SNRs facilitate higher feedback accuracy in terms of the average GCS. In the SNR regime higher than $20$ dB, the proposed IEFSF and all learning-based baselines achieve a performance similar to that in noise-free cases. It is worth noting that the non-learning-based baselines, i.e., the RVQ-based CSI feedback method and the Type I codebook-based CSI feedback method, are very sensitive to the noise since they directly feed back the indices of codewords to the BS, therefore they can not achieve a performance similar to that in noise-free cases even with a $30$ dB SNR. This further demonstrates the robustness of learning-based CSI feedback methods over conventional non-learning-based methods.
Figure 12: Average GCS of the proposed IEFSF and baselines versus SNRs under the indoor scenario.
Figure 13: Average GCS of the proposed IEFSF and baselines versus SNRs under the outdoor scenario.
@@
Finally, we compare the computational complexity of the proposed IEFSF with learning-based baselines. The computational complexity of all schemes with a fixed setting of feedback bits $B=256$ is shown in Table V. From Table V, the IEFSF and the SampleDL show computational efficiency over common DL-based CSI feedback schemes, i.e., CsiNet+, bi-ImCsiNet, and TransNet. This is because the SampleDL and the proposed IEFSF greatly reduce the CSI dimension for feedback before the CSI compression, such that the computational complexity can be significantly reduced. Further comparing the proposed IEFSF with the SampleDL, the proposed IEFSF has a significantly smaller time complexity than the SampleDL (i.e., $25.1$ times smaller) at the cost of a slightly larger space complexity (i.e., $1.3$ times larger than the SampleDL).
TABLE V: Computational complexity of the proposed IEFSF and baselines.
@@ VI Conclusion
VII Acknowledgement
This work was performed in part at SICC which is supported by SKL-IOTSC, University of Macau.